How many years will it take $3500 invested at 7% to earn $735?
Found 2 solutions by ewatrrr, Alan3354:
Answer by ewatrrr(24785)


Hi
A= p(1+r/n) ^nt (n=1 in this example: compounded annually
A = p(1+r)^t |A is p + I
log(A/p)/log(1+r) = t
log(4235/3500)/log(1.07) = t = 2.8 or 3 years

Answer by Alan3354(69443)

If it's simple interest (you didn't specify a compounding period),
$3500*0.07 = $245 interest per year
735/245 = 3 years